OpenStudy (anonymous):

What is the solution to this equation? 1/3x + 1/3(2x - 15) = 3 1/2 Solve for x

OpenStudy (whpalmer4):

\[\frac{1}{3}x + \frac{1}{3}(2x-15) = 3\frac{1}2\]Is that correct? I suggest starting by multiplying the entire equation by \(2*3=6\)